r2 - 17 Jan 2012 - 13:21:39 - KaushikDeYou are here: TWiki >  Admins Web > MinutesDataManageJan17



Minutes of the US ATLAS Data Management meeting, January 17, 2012
  • Previous meetings and background : IntegrationProgram
  • Coordinates: Tuesdays, Noon Central
    • (605) 475-6150, Access code: 735188; Dial *6 to mute/un-mute.


  • Meeting attendees: Armen, Dave, Rob, Bob, Wensheng, John, Saul, Hiro
  • Apologies:
  • Guests:

Topics for this week

  • Storage status, central deletion - Armen
    • No issues except adding more space to LOCALGROUPDISK in a few places, and cleaning up PRODDISK (lots of reco jobs recently)
    • Not much to delete recently
    • Shifters have very low threshold for reporting transfer errors to GGUS - 10 errors in 4 hours, being discussed
    • Some LFC errors again - at AGLT2 looks like ACL problems, NET2 not easy to understand
  • USERDISK cleanup - Hiro
    • Will send next round of emails this week - wait for ATLR outage effects to be minimized
  • Hot issues
  • AOB
    • DQ2 SS is in the process of moving to CERN (BNL has moved, next some Tier 2's)

-- KaushikDe - 17 Jan 2012

About This Site

Please note that this site is a content mirror of the BNL US ATLAS TWiki. To edit the content of this page, click the Edit this page button at the top of the page and log in with your US ATLAS computing account name and password.


Powered by TWiki
This site is powered by the TWiki collaboration platformCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback